Solve the differential equation d^2 y/dx^2 = a+bx+cx^2, given that dy/dx=0 and y= d, when x=0?
1
Expert's answer
2019-09-26T10:07:05-0400

"d^2 y\/dx^2 = a+bx+cx^2"


"d y/dx = ax+\frac{1}{2}bx^2+\frac{1}{3}cx^3+C_1"


"y(x) = \frac{1}{2}ax^2+\frac{1}{6}bx^3+\frac{1}{12}cx^4+C_1x+C_2"


"dy(0)/dx=0 \Rightarrow C_1=0"


"y(0) =d \Rightarrow C_2=d"

So, solution of differential equation


"y(x) = \frac{1}{2}ax^2+\frac{1}{6}bx^3+\frac{1}{12}cx^4+d"